An emf of 0.45 V is induced between the ends of a metal bar moving through a magnetic field of 0.22 T. What field strength would be needed to produce an emf of 1.5 V between the ends of the bar, assuming that all other factors remain the same?
Answer #1

as we know

e = BLv

considering all factors to be constant

e1/e2 = B1/B2

0.45/ 1.5 = 0.22 / B2

B2 = 0.73333 T

=======

Comment before rate in case any doubt, will reply for sure.. goodluck